MILLIMETER WAVE SPECTRA OF THE INTERNAL ROTATION EXCITED STATES OF (o)H_{2}-H_{2}O AND (o)H_{2}-D_{2}O

\underline{K. HARADA}, Y. IWASAKI, T. GIESEN, and K. TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka, 812-8581 JAPAN.

H_{2}-H_{2}O is a weakly bound complex and it has a various states according to the internal rotation for both H_{2} and H_{2}O moieties. In our previous study,^{a} we have reported the pure rotational transitions of the (o)H_{2} complex in the ground H_{2}O rotational state, 0_{00}(Σ), for both H_{2}-H_{2}O and H_{2}-D_{2}O, where (o)H_{2} (j_{H2} =1) is rotating perpendicular to the intermolecular axis to give the projection of j_{H2} to the axis k_{H2} to be zero (i.e. Σ state).

In the present study, we have observed the rotational transitions for the 0_{00} (Π) states in the millimeter-wave region up to 220 GHz, where the (o)H_{2} is rotating around the intermolecular axis to give the projection k_{H2} to be one (i.e. Π state). The center of mass bond lengths derived from the observed rotational constants for 0_{00}(Π) are longer by 5 % than those for 0_{00}(Σ), while force constants for the intermolecular stretching for 0_{00}(Π) derived from centrifugal distortion constants are smaller by 23 % than those for 0_{00}(Σ), suggesting the Π and Σ substates have quite different structures.

The recent theoretical calculation^{b} indicates that for 0_{00}(Σ), (o)H_{2} is bound to the oxygen site of H_{2}O, while for the 0_{00}(Π) state, (o)H_{2} to the hydrogen site of H_{2}O, and the 0_{00}(Σ) state is by 14 cm^{−1} more stable than the 0_{00}(Π) state. Observed molecular constants for 0_{00}(Σ) and (Π) are consistent with the structures given by the theoretical calculation.

We also observed the rotational spectrum in the 1_{01}(Σ) and (Π) states, where Σ and Π correspond to the rotation of H_{2}O perpendicular and parallel to the intermolecular axis and (o)H_{2} is calculated to be bound to the oxygen site of H_{2}O. The energy difference between the 1_{01}(Σ) and (Π) states will be discussed due to the Criolis interaction between these substates.

aC. J. Whitham, K. Tanaka, and K. Harada, The 56th OSU Symposium, RD08 (2001).

bAd. van der Avoid and D. J. Nesbit, J. Chem. Phys. 134, 044314 (2011).